# EvoC2Rust: A Skeleton-guided Framework for Project-Level C-to-Rust Translation

### v2

@@ Abstract.
Translating legacy C codebases to Rust is increasingly demanded for building safety-critical systems. While various approaches have emerged for this task, they face inherent trade-offs: rule-based methods often struggle to satisfy code safety and idiomaticity requirements, while LLM-based methods frequently fail to generate semantically equivalent Rust code, due to the heavy dependencies of modules across the entire codebase. Recent studies have revealed that both solutions are limited to small-scale programs. In this paper, we propose EvoC2Rust, an automated framework for converting complete C projects to equivalent Rust ones. EvoC2Rust employs a skeleton-guided translation strategy for project-level translation. The pipeline consists of three stages: 1) it first decomposes the C project into functional modules, employs a feature-mapping-enhanced LLM to transform definitions and macros, and generates type-checked function stubs, which form a compilable Rust skeleton; 2) it then incrementally translates functions, replacing the corresponding stub placeholders; 3) finally, it repairs compilation errors by integrating LLM and static analysis. Through evolutionary augmentation, EvoC2Rust combines the advantages of both rule-based and LLM-based solutions. Our evaluation on open-source benchmarks and six industrial projects demonstrates the superior performance of EvoC2Rust in project-level C-to-Rust translation. The results show that our approach outperforms the strongest LLM-based baseline by 17.24% in syntax accuracy and 14.32% in semantic accuracy, while also achieving a 43.59% higher code safety rate than the best rule-based tool.
@@ 1. Introduction
The translation of legacy C code to Rust is gaining significant attention in software engineering, spurred by critical memory safety concerns (Li et al., 2025; DARPA-SN-24-89, 2024). Such vulnerabilities constitute the most prevalent category of critical security defects in production C systems, accounting for approximately 70% of high-severity vulnerabilities in industry reports from Google and Microsoft (Lord, 2023). This has accelerated a paradigm shift toward compile-time safety guarantees through safe programming languages (Office of the National Cyber Director, 2024). In response, Rust has emerged as a leading alternative that offers memory safety by enforcing a strict ownership and borrowing model (Williams, 2024).
@@
Recent work proposes rule-based tools (c2r, 2025; Zhang et al., 2023; Hong and Ryu, 2023), which convert C code to Rust using primarily syntax-based transformations on a terse intermediate representation of compilation. Such translation inherently generates non-idiomatic Rust code with low-level constructs and may also contain semantic inaccuracies. (Pan et al., 2024). Another line of work proposes LLM-based techniques (Eniser et al., 2024; Yang et al., 2024b; Nitin and Ray, 2024) to utilize the code comprehension capability of code LLMs. Due to the lack of parallel C-Rust training data and significant linguistic differences, these approaches often fail to produce semantically equivalent Rust code (Ou et al., 2024).
While automated code translation has been extensively investigated (Rozière et al., 2020; Zhu et al., 2022; Yang et al., 2024a), the translation of C-to-Rust projects poses special challenges:
Challenge-1: Substantial linguistic discrepancies for ensuring safety. C and Rust differ fundamentally in their safety models. C permits loose type checking, unrestricted pointer arithmetic, and manual memory management, whereas Rust enforces strict compile-time type safety, ownership-based memory guarantees, and disciplined reference usage through its borrowing rules. Existing rule-based translators often compromise safety by relying on raw pointers, unsafe blocks, or external C functions (Pan et al., 2024). While LLM-based approaches can generate code with improved safety, they often lack sufficient context understanding (e.g., variable scope or pointer lifetimes), potentially leading to logical errors or memory safety violations in the translated Rust code (Sim et al., 2025). As a result, neither rule-based nor LLM-based methods alone reliably meet Rust’s stringent safety requirements.
Challenge-2: Project-level code dependence. Translating complete C projects poses greater challenges than translating isolated functions because the translated Rust project must preserve cross-module dependencies and consistent project-level structure (Shiraishi and Shinagawa, 2024; Shetty et al., 2024). Code LLMs often struggle with such large-scale context, leading to broken references, inconsistent APIs, or incorrect module boundaries (Pan et al., 2024). Consequently, existing approaches tend to perform well only on small-scale programs (e.g., $\leq$100 LoC) (Emre et al., 2021; Eniser et al., 2024), and fall short in real-world project settings.
@@
We evaluate EvoC2Rust using an open-source benchmark Vivo-Bench (Vivo, 2025) and a self-collected dataset of six industrial projects. Experimental results demonstrate that EvoC2Rust significantly outperforms existing baselines in project-level C-to-Rust translation. In comparison with LLM-based approaches, EvoC2Rust improves the compilation pass rate by 17.24–50.84% and the line acceptance rate by 14.32–44.79%, while also achieving marginally better code safety. When compared to rule-based methods, EvoC2Rust attains a 43.59–96.79% higher code safety rate, while maintaining competitive performance in terms of compilation and line acceptance rates. Furthermore, at the module level, EvoC2Rust achieves a 92.25% compilation success rate and an 89.53% test pass rate on industrial projects.
@@ 2.1. Project Skeleton Construction
Given a C project, EvoC2Rust first parses the source code using Tree-sitter (tre, 2025) to extract key syntactic elements and abstract them into a unified project metadata representation. This metadata captures structural information including filenames, include statements, macro definitions, type declarations, global identifiers, function signatures, and a mapping table between identifier declarations to source files. They are then systematically transformed into Rust equivalents via rule-based conversion. For instance, C include dependencies are transformed to Rust use imports, and identifier declarations are converted into Rust pub use statements (e.g., pub use {Rust_filename}::{identifier_name}) using the declaration-to-file table.
@@
Figure 2. An illustration of translating a simple C project to Rust.
@@ 3.1. Comparison Methods
  • C2Rust (c2r, 2025): a rule-based C-to-Rust translator that converts C code to Rust via AST analysis and manually defined transformation rules.
  • C2SaferRust (Nitin et al., 2025): a hybrid technique built on C2Rust, augmented with LLM-based post-processing and test-driven optimization to enhance the safety and idiomaticity of the translated Rust code. We follow the original configuration and use 5 optimization iterations.
  • Self-Repair (Khatry et al., 2025): a multi-agent self-repair framework for project-level C-to-Rust translation. It incrementally performs file translation and uses iterative compilation and testing feedback to correct errors automatically. We adopt the default setting of 3 self-repair rounds per task.
  • Tymcrat (Hong and Ryu, 2025b): a project-level translation method that improves translation quality through type migration. It generates multiple candidate Rust signatures and refines them iteratively using compiler feedback. We use the default configuration, which generates 4 candidate signatures per function.
@@ 3.2. Datasets
We evaluate EvoC2Rust using two benchmarks: the open-source Vivo-Bench (Vivo, 2025), and our newly introduced C2R-Bench, which comprises six industrial projects.
@@
Each project in both Vivo-Bench (Vivo, 2025) and C2R-Bench provides four artifacts: the original C source code, C test cases, the reference Rust translation, and corresponding Rust test cases. Table 3 summarizes the dataset statistics.
@@ 3.3. Evaluation Strategy and Metrics
  • Incremental Compilation Pass Rate (ICompRate): The proportion of translated modules that compile successfully when integrated step-wise into the project skeleton, measuring syntactic correctness  (Shiraishi and Shinagawa, 2024).
  • Line Acceptance Rate (AccRate) (Ziegler et al., 2022): This metric evaluates the fidelity of the initial translation by comparing it to the manually corrected version. Precision captures the percentage of correct lines in the initial output, while Recall measures their preservation in the final code.
  • Code Safety Rate (SafeRate) (Ling et al., 2022): The proportion of generated Rust code that is memory-safe, i.e., free of unsafe functions or blocks.
@@
  • Test Pass Rate (TestRate) (Khatry et al., 2025): The percentage of generated modules that pass all associated unit tests.
@@ 3.4. Implementation Details
C2Rust (c2r, 2025) 100 100 100 0
C2SaferRust (Nitin et al., 2025) 100 100 100 60.00
@@
Self-Repair (Khatry et al., 2025) 84.21 92.08 89.24 87.60
Tymcrat (Hong and Ryu, 2025b) 87.35 92.11 92.44 97.88
@@
Self-Repair (Khatry et al., 2025) 52.63 33.12 38.36 77.57
Tymcrat (Hong and Ryu, 2025b) 81.76 76.82 74.47 87.11
@@
C2Rust (c2r, 2025) 99.28 98.99 98.98 1.83
C2SaferRust (Nitin et al., 2025) 99.28 97.47 97.68 48.24
@@
Self-Repair (Khatry et al., 2025) 49.21 72.35 13.79 83.92
Tymcrat (Hong and Ryu, 2025b) 72.02 77.82 74.94 95.92
@@
Self-Repair (Khatry et al., 2025) 39.40 45.38 10.65 87.15
Tymcrat (Hong and Ryu, 2025b) 70.56 66.72 65.91 82.75
@@
We implement EvoC2Rust using DeepSeek-V3 (DeepSeek-AI, 2025) and Qwen3-32B (Alibaba, 2025) as the foundation models. Translation tasks use greedy decoding with max_tokens set to 4096. We leverage Tree-sitter v0.22.3 (tre, 2025) for C code parsing, the standard rustc (rus, 2025) for compilation, and the BGE-M3 model (of Artificial Intelligence, 2025) to generate embeddings and retrieve the top-10 most relevant patterns based on cosine similarity. Iteration limits are enforced to ensure practicality: 5 rounds for bracket repair and 3 rounds for LLM refinement.
Verified ground-truth implementations were established for the line acceptance rate metric as follows. The translation for each project was first executed by three senior engineers using the target methods. The generated Rust artifacts were then iteratively compiled, tested, and manually repaired with Claude Sonnet 4 (Anthropic, 2024) until all tests passed. A final code review by three authors confirmed the semantic equivalence of these implementations.
@@ 4.4. Scalability on Larger Projects (RQ4)
To evaluate the scalability and performance of our method on large-scale codebases, we conduct experiments on 10 larger projects (9,973 - 91,588 LoC) from RepoTransBench (Wang et al., 2024), a benchmark derived from highly-starred GitHub repositories. During module translation, we exclude functions that depend on third-party libraries, as our current implementation primarily supports dependencies on the ISO C standard library. Using DeepSeek-V3 as the backbone model, EvoC2Rust achieves acceptable performance, with an ICompRate of 75.61% and a SafeRate of 96.20%. We do not report line acceptance rates due to the prohibitive cost of manually creating verified Rust references for projects at this scale.
@@ 5. Limitations and Threats to Validity
Internal Validity. The main internal threat comes from correctness verification, which currently relies on predefined test cases. Adopting more robust validation techniques like fuzzing (Kadron et al., 2024) and self-debugging (Kang et al., 2025; Shi et al., 2024) could enhance test coverage and uncover subtle semantic errors. For line acceptance rate measurement, we use Claude-generated outputs verified by human experts as ground truth. While this manual validation ensures correctness beyond automated testing, it may not scale to larger datasets. Future work should explore semi-automated verification to balance accuracy and efficiency.
@@ 6.1. C-to-Rust Translation
Rule-based Methods rely on manually designed rules to transform C code into Rust. The C2Rust transpiler (c2r, 2025) exemplifies this approach by using Clang’s AST to generate Rust code, though the output relies heavily on unsafe blocks. Subsequent work has aimed to improve safety and idiomaticity: Emre et al. (Emre et al., 2021, 2023) used compiler feedback for pointer safety; Zhang et al. (Zhang et al., 2023) applied static ownership analysis; Ling et al. (Ling et al., 2022) and Hong et al. (Hong and Ryu, 2023, 2024a, 2024b, 2025a) focused on API safety and specific constructs; Larson et al. (Larsen, 2024) proposed pointer derivation graphs to translate unsafe Rust to safer Rust; Wu et al. (Wu and Demsky, 2025) retyped polymorphic C void pointers into uses of Rust generics. While effective, these methods require significant manual effort and often produce Rust code that is both unsafe and non-idiomatic.
LLM-based Methods leverage LLMs to generate idiomatic Rust code without relying on manually defined rules. Eniser et al. (Eniser et al., 2024) introduced FLOURINE, which uses fuzz testing to validate semantic equivalence between C and Rust code. Yang et al. (Yang et al., 2024b) adapted MSWasm-based testing for multi-language translation to Rust. Luo et al. (Luo et al., 2025) proposed IRENE, an LLM-based framework that integrates rule-augmented retrieval and structured summarization for function-level translation. Nitin et al. (Nitin and Ray, 2024) enhanced LLM translation by integrating static specifications and I/O tests. Xu et al. (Xu and Huang, 2025) optimized type migration by leveraging data flow graphs. Farrukh et al. (Farrukh et al., 2025) developed a multi-agent framework featuring iterative error correction. However, LLM-based translation suffers from lower accuracy due to the lack of parallel C-Rust training data (Ou et al., 2024).
Hybrid Methods seek to combine the strengths of both paradigms. For instance, C2SaferRust (Nitin et al., 2025) and PR2 (Gao et al., 2025) employ LLMs to refine rule-based translations, improving safety and idiomaticity while preserving functional equivalence through testing. SACTOR (Zhou et al., 2025) delegates different tasks to each paradigm, applying C2Rust for data types and an LLM for global variables and functions. Our method extends this direction with a two-stage framework: it first uses linguistic feature mappings (i.e., transformation rules) to guide LLM translation, then applies a compilation-driven repair mechanism combining LLM refinement with static analysis. It achieves an effective balance between translation accuracy and safety in automated C-to-Rust migration.

### v3

@@ Abstract.
Translating legacy C codebases to Rust is increasingly demanded for building safety-critical systems. While various approaches have emerged for this task, they face inherent trade-offs: rule-based methods often struggle to satisfy code safety and idiomaticity requirements, while LLM-based methods frequently fail to generate semantically equivalent Rust code, due to the heavy dependencies of modules across the entire codebase. In this paper, we propose EvoC2Rust, an automated framework for converting complete C projects to equivalent Rust ones. EvoC2Rust employs a skeleton-guided translation strategy for project-level translation. The pipeline consists of three stages: 1) it first decomposes the C project into functional modules, employs a feature-mapping-enhanced LLM to transform definitions and macros, and generates type-checked function stubs, which form a compilable Rust skeleton; 2) it then incrementally translates functions, replacing the corresponding stub placeholders; 3) finally, it repairs compilation errors by integrating LLM and static analysis. Through evolutionary augmentation, EvoC2Rust combines the advantages of both rule-based and LLM-based solutions. Our evaluation on open-source benchmarks and six industrial projects demonstrates the superior performance of EvoC2Rust in project-level C-to-Rust translation. The results show that our approach outperforms the strongest LLM-based baseline by 17.24% in syntax accuracy and 14.32% in semantic accuracy, while also achieving a 43.59% higher code safety rate than the best rule-based tool.
@@ 1. Introduction
The translation of legacy C code to Rust is gaining significant attention in software engineering, spurred by critical memory safety concerns (LiW25; DARPA). Such vulnerabilities constitute the most prevalent category of critical security defects in production C systems, accounting for approximately 70% of high-severity vulnerabilities in industry reports from Google and Microsoft (UrgentNeed). This has accelerated a paradigm shift toward compile-time safety guarantees through safe programming languages (WhiteHouse). In response, Rust has emerged as a leading alternative that offers memory safety by enforcing a strict ownership and borrowing model (Williams24a).
@@
Recent work proposes rule-based tools (c2rust; ZhangDYW23; HongR23), which convert C code to Rust using primarily syntax-based transformations on a terse intermediate representation of compilation. Such translation inherently generates non-idiomatic Rust code with low-level constructs and may also contain semantic inaccuracies. (LostinTranslation). Another line of work proposes LLM-based techniques (FLOURINE; LiYGCS24; vert; SpecTra) to utilize the code comprehension capability of code LLMs. Due to the lack of parallel C-Rust training data and significant linguistic differences, these approaches often fail to produce semantically equivalent Rust code (RustRepoTrans).
While automated code translation has been extensively investigated (JiaoYLQGS23; RoziereLCL20; Zhu0R22; Yang0YK0LHMJ024; TaoYGS24; dehghan2025translating), the translation of C-to-Rust projects poses special challenges:
Challenge-1: Substantial linguistic discrepancies for ensuring safety. C and Rust differ fundamentally in their safety models. C permits loose type checking, unrestricted pointer arithmetic, and manual memory management, whereas Rust enforces strict compile-time type safety, ownership-based memory guarantees, and disciplined reference usage through its borrowing rules. Existing rule-based translators often compromise safety by relying on raw pointers, unsafe blocks, or external C functions (LostinTranslation). While LLM-based approaches can generate code with improved safety, they often lack sufficient context understanding (e.g., variable scope or pointer lifetimes), potentially leading to logical errors or memory safety violations in the translated Rust code (LAC2R). As a result, neither rule-based nor LLM-based methods alone reliably meet Rust’s stringent safety requirements.
Challenge-2: Project-level code dependence. Translating complete C projects poses greater challenges than translating isolated functions because the translated Rust project must preserve cross-module dependencies and consistent project-level structure (projectc2rust; Syzygy). Code LLMs often struggle with such large-scale context (YunLGS24), leading to broken references, inconsistent APIs, or incorrect module boundaries (LostinTranslation; apirat). Consequently, existing approaches tend to perform well only on small-scale programs (e.g., $\leq$100 LoC) (EmreSDH21; FLOURINE), and fall short in real-world project settings.
@@
We evaluate EvoC2Rust using an open-source benchmark Vivo-Bench (vivo) and a self-collected dataset of six industrial projects. Experimental results demonstrate that EvoC2Rust significantly outperforms existing baselines in project-level C-to-Rust translation. In comparison with LLM-based approaches, EvoC2Rust improves the compilation pass rate by 17.24–50.84% and the line acceptance rate by 14.32–44.79%, while also achieving marginally better code safety. When compared to rule-based methods, EvoC2Rust attains a 43.59–96.79% higher code safety rate, while maintaining competitive performance in terms of compilation and line acceptance rates. Furthermore, at the module level, EvoC2Rust achieves a 92.25% compilation success rate and an 89.53% test pass rate on industrial projects.
@@ 2.1. Project Skeleton Construction
Figure 2. An illustration of translating a simple C project to Rust.
@@
Given a C project, EvoC2Rust first parses the source code using Tree-sitter (treesitter) to extract key syntactic elements and abstract them into a unified project metadata representation. This metadata captures structural information including filenames, include statements, macro definitions, type declarations, global identifiers, function signatures, and a mapping table between identifier declarations to source files. They are then systematically transformed into Rust equivalents via rule-based conversion. For instance, C include dependencies are transformed to Rust use imports, and identifier declarations are converted into Rust pub use statements (e.g., pub use {Rust_filename}::{identifier_name}) using the declaration-to-file table.
@@ 3.1. Comparison Methods
  • C2Rust (c2rust): a rule-based C-to-Rust translator that converts C code to Rust via AST analysis and manually defined transformation rules.
  • C2SaferRust (C2SaferRust): a hybrid technique built on C2Rust, augmented with LLM-based post-processing and test-driven optimization to enhance the safety and idiomaticity of the translated Rust code. We follow the original configuration and use 5 optimization iterations.
  • Self-Repair (CRUST-Bench): a multi-agent self-repair framework for project-level C-to-Rust translation. It incrementally performs file translation and uses iterative compilation and testing feedback to correct errors automatically. We adopt the default setting of 3 self-repair rounds per task.
  • Tymcrat (typemigrating): a project-level translation method that improves translation quality through type migration. It generates multiple candidate Rust signatures and refines them iteratively using compiler feedback. We use the default configuration, which generates 4 candidate signatures per function.
@@ 3.2. Datasets
We evaluate EvoC2Rust using two benchmarks: the open-source Vivo-Bench (vivo), and our newly introduced C2R-Bench, which comprises six industrial projects.
@@
Each project in both Vivo-Bench (vivo) and C2R-Bench provides four artifacts: the original C source code, C test cases, the reference Rust translation, and corresponding Rust test cases. Table 3 summarizes the dataset statistics.
@@ 3.3. Evaluation Strategy and Metrics
  • Incremental Compilation Pass Rate (ICompRate): The proportion of translated modules that compile successfully when integrated step-wise into the project skeleton, measuring syntactic correctness  (projectc2rust).
  • Line Acceptance Rate (AccRate) (0001KLRRSSA22): This metric evaluates the fidelity of the initial translation by comparing it to the manually corrected version. Precision captures the percentage of correct lines in the initial output, while Recall measures their preservation in the final code.
  • Code Safety Rate (SafeRate) (CRustS): The proportion of generated Rust code that is memory-safe, i.e., free of unsafe functions or blocks.
@@
  • Test Pass Rate (TestRate) (CRUST-Bench): The percentage of generated modules that pass all associated unit tests.
@@ 3.4. Implementation Details
C2Rust (c2rust) 100 100 100 0
C2SaferRust (C2SaferRust) 100 100 100 60.00
@@
Self-Repair (CRUST-Bench) 84.21 92.08 89.24 87.60
Tymcrat (typemigrating) 87.35 92.11 92.44 97.88
@@
Self-Repair (CRUST-Bench) 52.63 33.12 38.36 77.57
Tymcrat (typemigrating) 81.76 76.82 74.47 87.11
@@
C2Rust (c2rust) 99.28 98.99 98.98 1.83
C2SaferRust (C2SaferRust) 99.28 97.47 97.68 48.24
@@
Self-Repair (CRUST-Bench) 49.21 72.35 13.79 83.92
Tymcrat (typemigrating) 72.02 77.82 74.94 95.92
@@
Self-Repair (CRUST-Bench) 39.40 45.38 10.65 87.15
Tymcrat (typemigrating) 70.56 66.72 65.91 82.75
@@
We implement EvoC2Rust using DeepSeek-V3 (deepseekv3) and Qwen3-32B (Qwen3-32B) as the foundation models. Translation tasks use greedy decoding with max_tokens set to 4096. We leverage Tree-sitter v0.22.3 (treesitter) for C code parsing, the standard rustc (rustc) for compilation, and the BGE-M3 model (BGE-M3) to generate embeddings and retrieve the top-10 most relevant patterns based on cosine similarity. Iteration limits are enforced to ensure practicality: 5 rounds for bracket repair and 3 rounds for LLM refinement.
We established verified ground-truth implementations to evaluate the line acceptance rate metric. For each project, we first translated the code using the target methods. The resulting Rust artifacts were then iteratively compiled and tested. Any compilation or test failures were manually repaired by three senior engineers with assistance from Claude Sonnet 4 (claude4) until all tests passed.
@@ 4.4. Scalability on Larger Projects (RQ4)
To evaluate the scalability and performance of our method on large-scale codebases, we conduct experiments on 10 larger projects (9,973 - 91,588 LoC) from RepoTransBench (repotransbench), a benchmark derived from highly-starred GitHub repositories. During module translation, we exclude functions that depend on third-party libraries, as our current implementation primarily supports dependencies on the ISO C standard library. Using DeepSeek-V3 as the backbone model, EvoC2Rust achieves acceptable performance, with an ICompRate of 75.61% and a SafeRate of 96.20%. We do not report line acceptance rates due to the prohibitive cost of manually creating verified Rust references for projects at this scale.
@@ 5. Limitations and Threats to Validity
Internal Validity. The main internal threat comes from correctness verification, which currently relies on predefined test cases. Adopting more robust validation techniques like fuzzing (KadronNPBPS24) and self-debugging (KangCYL25; shi2024code) could enhance test coverage and uncover subtle semantic errors. For line acceptance rate measurement, we use Claude-generated outputs verified by human experts as ground truth. While this manual validation ensures correctness beyond automated testing, it may not scale to larger datasets. Future work should explore semi-automated verification to balance accuracy and efficiency.
@@ 6.1. C-to-Rust Translation
Rule-based Methods rely on manually designed rules to transform C code into Rust. The C2Rust transpiler (c2rust) exemplifies this approach by using Clang’s AST to generate Rust code, though the output relies heavily on unsafe blocks. Subsequent work has aimed to improve safety and idiomaticity: Emre et al. (EmreSDH21; EmreSDH23) used compiler feedback for pointer safety; Zhang et al. (ZhangDYW23) applied static ownership analysis; Ling et al. (CRustS) and Hong et al. (HongR23; HongR24a; HongR24b; Forcrat) focused on API safety and specific constructs; Larson et al. (Larsen24) proposed pointer derivation graphs to translate unsafe Rust to safer Rust; Wu et al. (WuD25) retyped polymorphic C void pointers into uses of Rust generics. While effective, these methods require significant manual effort and often produce Rust code that is both unsafe and non-idiomatic.
LLM-based Methods leverage LLMs to generate idiomatic Rust code without relying on manually defined rules. Eniser et al. (FLOURINE) introduced FLOURINE, which uses fuzz testing to validate semantic equivalence between C and Rust code. Yang et al. (vert) adapted MSWasm-based testing for multi-language translation to Rust. Luo et al. (IRENE) proposed IRENE, an LLM-based framework that integrates rule-augmented retrieval and structured summarization for function-level translation. Nitin et al. (SpecTra) enhanced LLM translation by integrating static specifications and I/O tests. Xu et al. (XuH25) optimized type migration by leveraging data flow graphs. Farrukh et al. (SafeTrans) developed a multi-agent framework featuring iterative error correction. However, LLM-based translation suffers from lower accuracy due to the lack of parallel C-Rust training data (RustRepoTrans).
Hybrid Methods seek to combine the strengths of both paradigms. For instance, C2SaferRust (C2SaferRust) and PR2 (PR2) employ LLMs to refine rule-based translations, improving safety and idiomaticity while preserving functional equivalence through testing. SACTOR (SACTOR) delegates different tasks to each paradigm, applying C2Rust for data types and an LLM for global variables and functions. Our method extends this direction with a two-stage framework: it first uses linguistic feature mappings (i.e., transformation rules) to guide LLM translation, then applies a compilation-driven repair mechanism combining LLM refinement with static analysis. It achieves an effective balance between translation accuracy and safety in automated C-to-Rust migration.